### v2

# Human Insights Driven Latent Space for Different Driving Perspectives: A Unified Encoder for Efficient Multi-Task Inference

Abstract
Autonomous driving systems require a comprehensive understanding of the environment, achieved by extracting visual features essential for perception, planning, and control. However, models trained solely on single-task objectives or generic datasets often lack the contextual information needed for robust performance in complex driving scenarios. In this work, we propose a unified encoder trained on multiple computer vision tasks crucial for urban driving, including depth, pose, and 3D scene flow estimation, as well as semantic, instance, panoptic, and motion segmentation. By integrating these diverse visual cues—similar to human perceptual mechanisms—the encoder captures rich features that enhance navigation-related predictions. We evaluate the model on steering estimation as a downstream task, leveraging its dense latent space. To ensure efficient multi-task learning, we introduce a multi-scale feature network for pose estimation and apply knowledge distillation from a multi-backbone teacher model. Our findings highlight two key findings: (1) the unified encoder achieves competitive performance across all visual perception tasks, demonstrating strong generalization capabilities; and (2) for steering estimation, the frozen unified encoder—leveraging dense latent representations—outperforms both its fine-tuned counterpart and the same frozen model pretrained on generic datasets like ImageNet. These results underline the significance of task-specific visual features and demonstrate the promise of multi-task learning in advancing autonomous driving systems. More details and the pretrained model are available at https://hi-computervision.github.io/uni-encoder/.
I Introduction
[…]
Figure 1: Our multi-task training strategy. $I_{s}$, $I_{t}$, $I_{1\dots 16}$ represent the source, target, and 16 sequential images, respectively. Their features, denoted as $f_{s}$, $f_{t}$, $f_{1\dots 16}$, are extracted (and concatenated when necessary) using our single encoder.
The advancement of self-driving cars has gained significant public attention in recent years, though the development of this technology began decades ago. Early innovations include vehicle-to-vehicle communication via radio waves in the 1920s [1] and electromagnetic guidance in the 1930s [2]. The primary goal of autonomous vehicles is to improve road safety and efficiency by minimizing human error, which causes more than 90% accidents in vehicles, while mechanical failures account for only 2% [3]. For fully autonomous driving cars, achieving human-level driving requires comprehensive environmental understanding, robust control, and reliable real-time decision-making.
A fully autonomous driving system usually extract a wide range of visual features—such as depth, motion, and segmentation—to support essential tasks like object detection, path planning, and behavior prediction. Beyond these explicit outputs, intermediate visual features (e.g., extracted by CNNs) that capture spatial relationships, temporal consistency, and scene dynamics are crucial for building a coherent understanding of the environment. These features enable downstream tasks, including obstacle avoidance, adaptive control, and trajectory prediction [4, 5, 6, 7]. Additionally, they play an essential role in interactive human-machine interfaces, which help explain how the system perceives and interprets its surroundings.
However, most learning-based models still focus on single-task objectives, such as steering estimation, using CNNs trained with RGB data to directly minimize the steering error [8, 9, 10, 11, 12]. While these approaches can capture relevant information, their reliance on narrow, task-specific datasets limits their capabilities. For example, Capito et al. [13] demonstrated that incorporating optical flow alongside RGB inputs significantly improved steering performance. This suggests that richer visual cues, beyond simple RGB data, can enhance both steering predictions and the overall contextual understanding of the driving scene.
In this paper, we propose a novel approach that integrates human-like perceptual information into a single encoder using a multi-task training strategy. We focus on computer vision tasks essential for urban navigation, including depth, pose, and 3D scene flow estimation, as well as semantic, instance, panoptic, and motion segmentation (see Figure 1). Our method provides several key advantages. First, after the initial training, the model can generate multiple outputs with a single pass through the unified encoder, reducing inference time while delivering diverse visual outputs. Additionally, our experimental results demonstrate that this approach achieves competitive performance compared to state-of-the-art methods across all individual tasks, ensuring the reliability of the extracted features. Finally, the encoder can be frozen and extended with a prediction head for steering estimation, enabling to assess the relevance of the learned features for this navigation task.
Training this model presents challenges in ensuring it performs well across all tasks. Naively training a shared encoder for depth and pose estimation led to drop in depth accuracy due to inaccurate pose estimation. Moreover, the lack of a comprehensive dataset with labeled data for all tasks led us to use a mixed training approach that combines supervised and self-supervised learning. However, this approach introduced its own issue: tasks with stronger training signals dominated the gradients in the encoder, overshadowing others. For instance, when training segmentation tasks (supervised learning) alongside 3D scene flow and motion mask tasks (self-supervised learning), the latter two struggled to learn effectively. Since these self-supervised tasks are critical for depth estimation in dynamic scenes, this training dominance destabilized the entire self-supervised training process. To this end, the contributions of this paper are:
  • A multi-scale pose decoder: We design a pose decoder that leverages multi-scale features from the shared encoder, improving depth estimation in dynamic scenes.
  • Knowledge distillation for stability: We apply knowledge distillation from a multi-encoder teacher model to guide the learning of self-supervised tasks and prevent gradient imbalances.
  • Comprehensive evaluation of shared features: We conduct a comprehensive evaluation of the unified encoder’s performance across all integrated tasks. Our results demonstrate that, for steering angle estimation, the frozen unified encoder outperforms its fine-tuned counterpart and the same architecture pretrained solely on generic datasets such as ImageNet.
II Related works
II-A Image segmentation
Image segmentation involves separating an image into segments, grouping pixels by specific criteria. Semantic segmentation assigns pixels to broad classes like roads and buildings, while instance segmentation focuses on distinct objects like cars and people [14]. Traditionally, these tasks were handled separately. To integrate these approaches, Kirillov et al. [15] introduced panoptic segmentation, which organizes pixels into amorphous background regions (”stuff”) and distinct objects (”things”). However, this led to an additional specialized task rather than unifying these two, with performance falling short of state-of-the-art results in dedicated tasks.
Recent methods have proposed unified models for all three tasks showing high performance but still requiring separate training for each. In a significant advancement, Jain et al. introduced OneFormer [16], a model that, given an image and a text prompt specifying the task, produces the corresponding segmentation output. Evaluated on several public datasets, OneFormer set new state-of-the-art benchmarks for all three tasks with a single, jointly trained model. In this paper, we use OneFormer decoders to generate the three mentioned segmentation outcomes.
[…]
Figure 2: Simplified architecture of our model: (a) Depth network using target image features $f_{t}$ to output depth $\mathbf{d}_{t}$, (b) Multi-scale pose network using source and target image features $f_{s},f_{t}$ to output relative pose $\mathbf{T}_{t\rightarrow s}$, (c) 3D Scene Flow $\mathbf{F}_{C}$ and Motion mask $\mathbf{M}$ networks using RGB images and features $f_{s},f_{t}$, (d) Segmentation network outputting panoptic, instance, and semantic segmentations, and (e) Loss computation $L_{ssup}$ for joint training of depth, pose, 3D scene flow, and motion mask segmentation. We denote rigid flow $\mathbf{F}_{R}$, independent flow $\mathbf{F}_{I}$, final flow, and sampled target image $\hat{\mathbf{I}}_{t}$.
II-B Monocular Depth & Pose Estimation
Monocular depth estimation is a task of predicting the depth of a scene from a single 2D image. Unlike stereoscopic methods, which use different viewpoints as input, monocular depth estimation must infer depth from just one, making it particularly challenging. Traditional methods mostly based on hand-crafted features, which can lead to inaccuracies in complex scenarios [17]. In a first attempt using deep learning, Eigen et al. (2014) introduced a CNN-based model to predict depth maps directly from single images, achieving superior results. However, this method requires ground truth for depth, typically obtained with expensive hardware like LiDAR, limiting its practicality. To overcome this, recent research has shifted to unsupervised methods that use the inherent structure of unlabeled images. Notably, Zhou et al. [18] and Godard et al. [19] proposed unsupervised methods which employ video sequences to simultaneously learn depth and camera motion, reducing the dependency on labeled data and adapting better to dynamic scenes.
However, unsupervised methods often consider the assumption of a static world, which is not true in most real scenarios. To this end, recent advancements have intergrated flow (e.g., 2D or 3D) and motion segmentation in addition to depth and pose estimation, enabling better handling of dynamic objects in the scene [20].
II-C Scene Flow and Motion Segmentation
Scene flow estimation is similar to depth estimation but in addition, it can capture 3D motion between consecutive images, while motion segmentation identifies dynamic objects. These tasks, when jointly trained with depth and pose estimation, improve depth accuracy [21, 22]. Jiao et al. ’s EffiScene network [23] trains on these tasks using stereo images, leveraging the shared geometric structure of scene depth and object movement. Recently, Sun et al. [20] propose DynamoDepth framework that further improves flexibility with the capacity of training on only video frames.
Building on DynamoDepth, we propose a single encoder model that requires only two images to estimate scene flow and motion segmentation, reducing complexity compared to the three-image requirement in the original method.
II-D Steering estimation
The two main research directions for steering estimation are model-based and model-free. Model-based methods rely on vehicle dynamics models, while model-free methods leverage data-driven techniques like deep learning. This paper focuses only on the model-free approach. Bojarski et al. [10] introduced an end-to-end steering estimation using CNN with a single RGB camera input. Capito et al. [13] improved the navigation capability by incorporating optical flow into the input. The importance of temporal information has also been highlighted in many works. Indeed, Eraqi et al. [24] employed LSTM networks to enhance steering control, while Xu et al. [25] combined fully convolutional networks with LSTM and semantic segmentation for improved road condition interpretation. Lechner et al. [4] proposed Neural Circuit Policies (NCPs), which provide interpretable decision maps from high-dimensional inputs. These studies indicate that integrating temporal information between frames can improve steering accuracy.
II-E Multi-Task Learning in Computer Vision
Recent works have explored multi-task learning (MTL) frameworks to simultaneously address multiple vision tasks, leveraging shared representations for improved performance. For instance, online knowledge distillation techniques have been proposed to mitigate negative transfer between tasks such as semantic segmentation and depth estimation, enhancing overall learning stability [26]. Joint-confidence-guided MTL frameworks have also been introduced for 3D reconstruction, combining depth prediction, semantic labeling, and surface normal estimation to improve feature fusion [27]. Moreover, collaborative MTL approaches have demonstrated effectiveness in handling object detection, segmentation, and tracking tasks by establishing associative connections among task heads [28].
Inspired by this synergy, we take a different perspective. Instead of demonstrating how MTL can enhance the performance of one or a few task through training, we hypothesize that the shared features learned from a diverse set of navigation tasks are sufficient not only for achieving good performance on each individual task but also for supporting a related navigation task—steering estimation—which is not present during training.
III Methods
Figure 1 illustrates our model architecture, which includes a Swin Tiny encoder [29] and six decoders for: panoptic, instance, and semantic segmentation, depth and pose estimation, 3D scene flow estimation, motion mask segmentation, and steering command prediction. Training was performed in two stages. In the first stage, we pretrained the encoder with five decoders (excluding the steering command decoder) to learn generalized features across tasks. In the second stage, we froze the encoder and added a prediction head for steering estimation to evaluate its performance in navigation.
Given the challenge of finding an annotated dataset for all the targeted tasks, we selected the CityScapes dataset [30] for training due to its high-quality ground truth segmentation and structured video sequences. This sequential format is particularly suitable for self-supervised training in depth, optical flow, and motion mask estimation without requiring explicit labels. However, we did not evaluate the steering prediction in CityScapes. Instead, this evaluation is performed on the MIT dataset [4] after fine-tuning a prediction head (with the frozen encoder), which provides the appropriate annotations for this task. CityScapes dataset serves as a crucial foundation for training the unified encoder, allowing it to capture rich visual and motion cues that contribute to downstream tasks (e.g., steering estimation).
III-A Panoptic, Instance, Semantic Segmentations
For these three supervised segmentation tasks, we employ the OneFormer approach. This is a task-conditioned universal image segmentation model that achieves state-of-the-art performance across all segmentation tasks with a unified model and architecture. The OneFormer framework adopts a task-conditioned joint training strategy that allows simultaneous training on all segmentation tasks using a single model, reducing resource requirements significantly.
Figure 2d provides a simplified illustration of the OneFormer architecture. This model consists of an encoder and two distinct decoders. The first decoder processes image features to generate all segment-related features. The output from this decoder, along with an embedding of a text prompt defining the task (e.g., ”The task is {panoptic, instance, semantic}”), is then passed to the second decoder. The second decoder uses this information to produce the desired segmentation type. The supervised loss function combines cross-entropy loss for classification, binary cross-entropy for mask predictions, and Dice loss for accurate mask boundary predictions. We trained this model end-to-end on panoptic annotations, from which semantic and instance labels are derived. The overall loss function $\mathbf{L}_{\text{sup}}$ integrates multiple components to ensure task-specific accuracy:
$\mathbf{L}_{\text{sup}}=\lambda_{\text{cls}}L_{\text{cls}}+\lambda_{\text{bce}}L_{\text{bce}}+\lambda_{\text{dice}}L_{\text{dice}}+\lambda_{\text{contrast}}L_{\text{contrast}},$ (1)
where:
  • $L_{\text{cls}}$ is the cross-entropy loss for mask (i.e., region) classification accuracy.
  • $L_{\text{bce}}$ is the binary cross-entropy loss applied to mask.
  • $L_{\text{dice}}$ is the Dice loss, helps to improve mask boundaries.
  • $L_{\text{contrast}}$ is the contrastive loss between object and text queries to make sure the text query is taken into account.
III-B Depth, Pose, 3D Flow Estimation, Motion Mask
For depth estimation, we employ the depth decoder architecture from [31]. In addition, we introduce a novel pose decoder, which is discussed in detail in the next section. For 3D scene flow estimation and motion mask segmentation, we build upon the DynamoDepth approach [20], enhancing it with convolutional blocks including additional non-linearity to increase the model’s representational capacity. Concretely, instead of considering the 3D scene flow as a linear combination of the encoder features $f$ and the pose $\mathbf{T}_{t\rightarrow s}$, we consider it as a non linear combination $\mathbf{F}_{i}=\mathbf{U}(\mathbf{F}_{i+1})+\textit{ConvELU}([a_{i},\textit{Conv}(a_{i}))]$ where $a_{i}=\textit{Conv}([\mathbf{U}(\mathbf{F}_{i+1}),f_{i}]])$, $\mathbf{F}_{5}=\mathbf{T}_{t\rightarrow s},\text{ and }\mathbf{F}\text{ and }\mathbf{U}$ are respectively flow and upscaling operation. By doing this, we increase the complexity of the decoders, which can help the unified encoder to learn simpler and more generalizable features that can effectively support downstream tasks. Figure 2a, b, and c provides simplified illustrations of our networks.
Starting with source $\mathbf{I}_{s}$ and target frames $\mathbf{I}_{t}$, we estimate depth $\mathbf{d}_{t}$ and relative pose $\mathbf{T}_{t\rightarrow s}$. Using $\mathbf{T}_{t\rightarrow s}$, we calculate 3D rigid flow $\mathbf{F}_{R}$. The source and target features are concatenated and passed through the 3D scene flow and motion mask segmentation decoders along with $\mathbf{T}_{t\rightarrow s}$, producing the complete flow $\mathbf{F}_{C}$ and binary motion mask $\mathbf{M}$. The independent flow $\mathbf{F}_{I}$ is computed as $\mathbf{F}_{I}=\mathbf{M}\times(\mathbf{F}_{C}-\mathbf{F}_{R})$. The final flow $\mathbf{F}_{F}=\mathbf{F}_{R}+\mathbf{F}_{I}$ is used to estimate the target frame $\hat{\mathbf{I}}_{t}$ from $\mathbf{I}_{s}$. The self-supervision loss $\mathbf{L}_{\text{ssup}}$ is:
$\mathbf{L}_{\text{ssup}}=L_{\text{recon}}+L_{s}+\lambda_{c}L_{c}+\lambda_{m}L_{m}+\lambda_{g}L_{g},$ (2)
where:
  • $L_{\text{recon}}$ is the reconstruction loss that combines SSIM and L1 norms weighted by $\alpha$.
  • $L_{s}$ is the smoothness loss [19] which is a weighted sum of edge-aware smoothness losses for the inverse depth, the $\mathbf{F}_{C}$ and $\mathbf{M}$.
  • $L_{c}$ is the motion consistency loss that penalize flow discrepancy $\mathbf{F}_{D}=\|\mathbf{F}_{C}-\mathbf{F}_{R}\|_{1}$ for static pixels.
  • $L_{m}$ is the motion sparsity loss that penalizes the mask M with low flow discrepancies using cross-entropy to favor zero motion mask.
  • $L_{g}$ is the above-ground loss that penalizes points projected below the ground plane using RANSAC.
Figure 3: Qualitative results. Left to right: Input, panoptic, instance, semantic output, depth, motion mask, independant flow.
Figure 2e illustrates the data flow for computing the self-supervised loss described above.
III-C Towards a Unified Multi-Task Encoder
Unlike the DynamoDepth method, which employs three distinct encoders—one each for depth estimation, pose estimation, and 3D scene flow with motion mask segmentation—we propose a single, unified encoder capable of handling all these tasks, along with three additional segmentation tasks: panoptic, instance, and semantic segmentation. This approach can enhance compactness and efficiency but also introduces several challenges.
The foremost challenge is the shared use of the encoder for both depth and pose estimation tasks, which can compromise depth estimation accuracy due to suboptimal relative pose estimation, as noted in [19]. In the shared encoder approach described in [19], relative pose estimation relies on the concatenated lowest-resolution features from the source and target images. However, these low resolution features can lead to ambiguity in pose estimation between the two images.
To overcome this limitation, we propose a multi-scale pose decoder that leverages multi-scale features from both the source and target images to improve the accuracy of relative pose estimation. Specifically, the images are processed through the shared encoder to extract multi-scale features. These features are then concatenated (as illustrated in Figure 2b), and convolutional blocks with skip connections are applied to each set of concatenated features, from high to low resolution. Given an input feature $f$, the output of the convolutional block with skip connections is: $\text{out}=\textit{ReLU}(\textit{ConvBNReLU}(\textit{ConvBN}(f))+\textit{Shortcut}(f))$ where Conv, BN, ReLU refers to the convolution, batch norm layers and ReLU activation. We apply two consecutive blocks to every input. Higher-resolution outputs are downsampled and concatenated with the lower resolution inputs before being passed through the next two block. Based on the obtained lowest resolution features, the final relative pose estimation is performed using the same pose prediction head as in [19].
A secondary challenge arises when training supervised and self-supervised tasks simultaneously. Some self-supervised tasks, such as 3D scene flow estimation and motion mask segmentation, struggle to converge when trained alongside the three supervised segmentation tasks using a single shared encoder. It is possible that the gradients from the supervised tasks are sufficiently strong to overshadow those from the self-supervised tasks, hindering their convergence.
To address this issue, we adopt a knowledge distillation strategy. We first train a teacher model consisting of three encoders and five decoders similar to [20]: one Swin encoder for depth estimation and the three supervised segmentation tasks, one ResNet18 encoder for pose estimation, and one ResNet18 encoder for 3D scene flow and motion mask segmentation. We then use the encoder-decoder pair responsible for 3D scene flow and motion mask segmentation in the teacher model to supervise these two tasks in the unified encoder model. The associated distillation loss is defined as:
$\mathbf{L}_{\text{distil}}=\beta_{1}\cdot\|\mathbf{F}_{C_{\text{teacher}}}-\mathbf{F}_{C_{\text{student}}}\|_{1}+\beta_{2}\cdot\|\mathbf{M}_{\text{teacher}}-\mathbf{M}_{\text{student}}\|_{1}$ (3)
Finally, we train the entire framework using a weighted sum of the supervised, self-supervised, and distillation losses:
$\textbf{L}_{total}=\lambda_{1}\cdot\textbf{L}_{sup}+\lambda_{2}\cdot\textbf{L}_{ssup}+\lambda_{3}\cdot\mathbf{L}_{\text{distil}}$ (4)
III-D Efficient Steering Estimation from Dense Latent Space
For the steering command, we used the pre-trained encoder, which was frozen, followed by an attentive pooling mechanism [32] to process the steering input. The encoder’s outputs were fed into the attentive pooler, which takes a sequence of 16 images to compute the steering angle. This approach effectively used the encoder’s latent representations while applying steering-specific attention through the pooling mechanism. During training, we optimized the loss $L_{\text{pred}}$, which is defined as follows:
$L_{\text{pred}}={\sum_{i}w_{i}(\hat{\mathbf{y}}^{(i)}-\mathbf{y}^{(i)})^{2}}\big{/}{\sum_{i}w_{i}},$ (5)
where $w_{i}=\exp\left(\lambda\cdot|\mathbf{y}^{(i)}|\right)$, with $\lambda$ representing the factor modulating the influence of the steering command’s magnitude, $|\mathbf{y}^{(i)}|$, on the loss. To evaluate the model’s performance, we employed the mean squared error (MSE) metric.
IV Experimental results
IV-A Training Setup
We use the KITTI Eigen split for initial ablation (trained only with the depth and the pose decoders) and CityScapes for the remaining experiments, with image resolutions of $192\times 640$ and $192\times 512$, respectively. Images are preprocessed into triples using scripts from [18]. For CityScapes, the lower 25% of images is cropped to exclude the front car [33]. The entire CityScapes dataset is used for supervised segmentation tasks. Pretraining is conducted on a single NVIDIA RTX A5000 with a batch size of 6 (3 images for the supervised tasks, i.e., segmentations, and 3 triples for the self-supervised tasks) over four steps: (1) training the shared encoder with depth, pose, and segmentation decoders for 60,000 steps, (2) training the 3D scene flow decoder for 40,000 steps with other components frozen, (3) training the pose decoder, 3D scene flow decoder, and motion mask segmentation decoder for 40,000 steps with other components frozen, and (4) training the whole network for 250,000 steps. After pretraining, we follow the tenfold cross-validation procedure by Lechner et al. [4] to fine-tune the model on the steering angle estimation task. The values of hyperparameters are: $\lambda_{\text{cls}}=2,\lambda_{\text{bce}}=\lambda_{\text{dice}}=\lambda_{\text{c}}=5,\lambda_{\text{contrast}}=0.5,\lambda_{\text{m}}=\lambda_{\text{g}}=\beta_{1}=0.1,\beta_{2}=1e-3,\lambda_{\text{1}}=\lambda_{\text{2}}=\lambda_{\text{3}}=1$.
IV-B Ablation Studies
Multi-scale pose decoder. To evaluate the effectiveness of our proposed multi-scale pose decoder, we conducted depth estimation experiments using the KITTI Eigen split dataset. This dataset was selected because accurate depth estimation can be achieved with just a depth and a pose network, allowing us to more clearly isolate and assess the contribution of our multi-scale pose decoder. Table I-1 presents the depth estimation results on the KITTI dataset. Our findings reveal that a naive approach of using a shared encoder—by concatenating the lowest-level features of source and target images and passing them through a pose prediction head—results in a decline in depth estimation performance. This trend is evident in both the Monodepth2 model [19] and our Swin encoder. In contrast, using our multi-scale pose decoder maintains depth performance comparable to using a separate ResNet18 encoder specifically for pose estimation.
Model Error (↓) Acc. (↑)
MS FM KD $Abs_{rel}$ $rmse_{log}$ $\delta<1.25$
$MD2_{sep}$ 0.115 0.193 0.877
$MD2_{sh}$ 0.125 0.201 0.857
$\text{Our}_{sep}$ 0.108 0.183 0.884
$\text{Our}_{sh}$ 0.117 0.190 0.872
$\text{Our}_{sh}$ 0.109 0.184 0.884
$\text{Our}_{sep}$ 0.103 0.157 0.885
$\text{Our}_{sh}$ 0.126 0.182 0.850
$\text{Our}_{sh}$ 0.134 0.183 0.833
$\text{Our}_{sh}$ 0.106 0.158 0.888
TABLE I: 1) (Upper part) Ablation on multi-scale pose decoder using KITTI dataset. 2) (Lower part) Analysis on knowledge distilation on CityScapes. MS, FM, and KD refers multi-scale pose decoder, 3D scene Flow & Motion mask, and Knowledge Distilation. sep and sh refer to the separate and shared version. For the separate encoder, a ResNet18 was employed similar to [19].
$PQ$ (↑) $AP$ (↑) $IoU$ (↑)
OneFormer [16] 55.8 28.4 74.3
Our multi-task model 56.0 28.6 74.2
TABLE II: Ablation study on panoptic, instance, and semantic segmentation tasks.
Method IM Sem. #$\#f$ D Error metric ($\downarrow$) Accuracy metric ($\uparrow$)
$Abs_{rel}$ $Sq_{rel}$ $RMSE$ $RMSE_{log}$ $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
Monodepth2 [19] 1 K 0.115 0.903 4.863 0.193 0.877 0.959 0.981
LiteMono [34] 1 K 0.101 0.729 4.454 0.178 0.897 0.965 0.983
Struct2Depth [35] 1 K 0.141 1.026 5.290 0.215 0.816 0.945 0.979
SGDepth [36] 1 K 0.113 0.835 4.693 0.191 0.879 0.961 0.981
Lee et al. [37] 1 K 0.124 0.886 5.061 0.206 0.844 0.948 0.979
RM-Depth [38] 1 K 0.107 0.687 4.476 0.181 0.883 0.964 0.984
Dynamo-Depth [20] 1 K 0.112 0.758 4.505 0.183 0.873 0.959 0.984
Ours (wo 3D scene flow) 1 K 0.109 0.818 4.654 0.184 0.884 0.963 0.983
Struct2Depth [35] 1 CS 0.145 1.737 7.280 0.205 0.813 0.942 0.978
Gordon et al. [39] 1 CS 0.127 1.330 6.960 0.195 0.830 0.947 0.981
Li et al. [40] 1 CS 0.119 1.290 6.980 0.190 0.846 0.952 0.982
Lee et al. [37] 1 CS 0.111 1.158 6.437 0.182 0.868 0.961 0.983
RM-Depth [38] 1 CS 0.100 0.839 5.774 0.154 0.895 0.976 0.993
Zhong et al. [41] 2 CS 0.098 0.946 5.553 0.148 0.908 0.977 0.992
DynamicDepth [42] 2 CS 0.103 1.000 5.867 0.157 0.895 0.974 0.991
ManyDepth [33] 2 CS 0.114 1.193 6.223 0.170 0.875 0.967 0.989
Ours 1 CS 0.106 1.033 5.913 0.158 0.888 0.974 0.982
TABLE III: Depth evaluation on the KITTI (K), and CityScapes (CS) Dataset. IM and Sem. stand for independent motion and semantics. #$\#f$ indicates the number of frames during inference. D is the used datatset for training and evaluation. Bold is best. Underline is ours.
Knowledge distillation for 3D flow & motion mask. We evaluate the effectiveness of knowledge distillation in training our unified encoder on the CityScapes dataset, which includes many dynamic objects, making 3D scene flow and motion mask networks crucial. Table I-2 shows the depth error and accuracy for different models. First, using a shared encoder with our multi-scale pose decoder achieves good performance compared to using separate encoders for different tasks. Second, training the shared encoder without knowledge distillation leads to suboptimal performance, even worse than the shared encoder model that doesn’t consider 3D scene flow and motion masks. However, incorporating knowledge distillation improves the shared encoder’s performance, making its performance inline with models with separate encoders. This highlights the effectiveness of our strategy.
Panoptic, instance, and semantic segmentations. To make a fair comparison, we trained OneFormer using the same batch size as our multi-task model. Table II shows the results for Panoptic Quality (PQ), Average Precision (AP), and Intersection over Union (IoU) respectively for panoptic, instance, and semantic segmentation tasks. We can observe that the segmentation performance of our multi-task model closely matches that of the OneFormer model, confirming the state-of-the-art capability of our approach.
IV-C Comparison with State-of-the-Art Methods
Since our multi-task model aligns closely with the state-of-the-art OneFormer model for segmentation tasks, we only focus on comparing depth estimation with other leading methods. Table III presents depth estimation results on the KITTI and CityScapes datasets.
On KITTI, due to the lack of labeled segmentation data, our model is trained without supervision but still surpasses the average performance of state-of-the-art methods. Notably, it outperforms several established methods (e.g., Struct2Depth) across all metrics, even without independent motion and segmentation support. Similarly, on CityScapes, our model shows competitive results, performing better than some methods and approaching the best results for each metric. It is important to highlight that most other methods rely on off-the-shelf segmentation models or use multiple encoders for pose and/or motion estimation.
Finally, Figure 3 shows qualitative results of our multi-task model. The panoptic, instance, and semantic segmentation demonstrate high quality, consistent with the strong quantitative results. Additionally, the quality of the estimated flow and motion mask demonstrate the contribution of these outputs to improved depth estimation compared to the model without the 3D scene flow and motion mask network.
Model Training error Test error
CNN-GRU 64 units [4] $1.25\pm 1.02$ $5.06\pm 6.64$
CNN-LSTM 64 units [4] $\mathbf{0.19}\pm 0.05$ $\mathbf{3.17}\pm 3.85$
VAE-LSTM 64 units [7] $0.54\pm 0.26$ $4.70\pm 4.80$
VAE-LSTM 19 units [7] $0.60\pm 0.30$ $6.75\pm 8.33$
(1) Swin-AttnPool (ImageNet pretrained) $2.91\pm 2.23$ $11.03\pm 10.03$
(2) Swin-AttnPool (Encoder unfrozen) $9.62\pm 3.24$ $16.46\pm 11.13$
(3) Swin-AttnPool (Encoder frozen) $1.64\pm 1.63$ $5.41\pm 6.06$
TABLE IV: Results from the passive lane-keeping evaluation across tenfold cross-testing. Bold is best. Underline is ours.
IV-D Dense Latent Space to Steering: Evaluation
Table IV summarizes the performance of various models for steering prediction, including our proposed Swin-AttnPool approach and several existing CNN-based [4] and VAE-based models [7]. To evaluate the effectiveness of the features learned through our training strategy, we present three variants of the same encoder architecture:
  • ImageNet-pretrained (frozen): The encoder is pretrained on ImageNet, frozen, and fine-tuned on the steering estimation task.
  • Our pretrained (unfrozen): The encoder is initialized with our pretrained weights and unfrozen during fine-tuning.
  • Our pretrained (frozen): The encoder is initialized with our pretrained weights and remains frozen during fine-tuning.
The results reveal several key insights. First, our frozen encoder (variant 3) outperforms the frozen encoder pretrained on ImageNet (variant 1), indicating that the visual features learned through our multi-task training are more relevant for navigation tasks. Second, the frozen encoder (variant 3) also surpasses its unfrozen counterpart (variant 2), suggesting that exclusively fine-tuning for the steering task may overlook valuable features learned from other navigation-related tasks, which can positively contribute to steering performance.
Finally, although Table IV shows that our training error (variant 3) is relatively high, the test error remains competitive compared to the average performance of all methods. Indeed, our frozen encoder outperforms the VAE-LSTM (19 units) model and performs comparably to the CNN-GRU (64 units) approach. The relatively high error could be due to architectural aspects, either within the encoder or the prediction head, that may not be fully optimized for the steering estimation task.
V Discussion
Training a unified encoder for multiple tasks introduces significant optimization challenges, particularly in balancing supervised and self-supervised losses. One notable issue was the dominance of segmentation loss over motion-based tasks, resulting in imbalanced feature learning. To address this, we introduced a knowledge distillation approach from a multi-encoder teacher network, which guided the learning of motion segmentation and scene flow, ensuring stable convergence. Additionally, the proposed multi-scale pose decoder enhanced depth estimation, mitigating inaccuracies in dynamic scenes. These improvements preserved the advantages of multi-task learning while addressing convergence issues.
Our approach represents one of the first attempts to integrate human-like perceptual information into a shared encoder, achieving performance comparable to multi-encoder methods in the literature. We argue that accurate autonomous driving requires human-like perception and decision-making. By leveraging a dense feature space enriched with human-understandable representations, the model captures richer, more contextual information about the driving scene. This leads to more robust and interpretable decisions in complex real-world scenarios.
Although the strategy shows promising results, higher numerical errors were observed in certain evaluations. These discrepancies could stem from two factors: (1) the used encoder architecture may not be fully optimized for the steering estimation task, and (2) the prediction head was not extensively analyzed, as we only employed an attention pooler among many possible alternatives. Future work should explore various prediction head architectures and refine the model to align more closely with task-specific requirements, potentially reducing these errors while preserving the benefits of multi-task learning.
Additionally, while our approach demonstrates the potential of multi-task learning, it introduces increased computational complexity due to the simultaneous training of multiple tasks. Future research should explore more efficient training strategies to optimize performance and scalability.
VI Conclusion
In this work, we presented a unified encoder architecture for autonomous driving, achieved through multi-task learning that incorporates human-like visual perception necessary for navigation. By integrating knowledge from critical computer vision tasks—including depth, pose, 3D scene flow estimation, and various segmentation tasks—into a single encoder, our approach enables efficient and compact multi-task inference.
We addressed key challenges associated with multi-task training to ensure the effectiveness of each individual task. First, we introduced a novel multi-scale pose decoder, which enhances relative pose estimation between frames and improves depth performance, especially in dynamic scenes such as those in the KITTI dataset. Second, we employed knowledge distillation from a multi-encoder teacher model pretrained on the same tasks to stabilize the training process. Our experimental results demonstrate that the unified encoder achieves performance on par with state-of-the-art methods for individual tasks. Moreover, when directly leveraging the latent space from the unified encoder, the architecture is capable of coherent steering angle estimation, benefiting from the diverse encoded visual information. This highlights the potential of multi-task learning to advance robust and interpretable autonomous navigation systems.